Instrument Log

SORTI 01       C1                  Lamont, OK Optical Trailer

Date

Operator

Comments

1/14/13

CAM

After reinstalling the SWS, noticed the dry air system was not providing air.  T/S found the copper air line between the compressor and the building was plugged with ice.  Disconnected copper line at building entrance, thawed the ice, and cleared water from line.  Reconnected line and confirmed air flow.  RTS at 2030 GMT.   SAS-He, RSS and SORTI were also affected by lack of air.

9/26/11

PAD

Purge Gas Generator located in Optical trailer has been repaired.  Unit needs to be RTS.  Installed Whatman FT-IR Model 75-45-12VDC, SN 7545-0236 Purge Gas Generator and connected inlet line.  Started unit and opened inlet valve to allow air to flow through system.  Vendor recommends unit allow to run 24 hours before connecting outlet line.  Will connect outlet line tomorrow.  Instruments affected are SWS, RSS, and SORTI.  RTS at 2120 GMT.

8/4/2011

PAD

Purge Gas Generator located in Optical trailer that services RSS, SWS, and SORTI is not operating correctly.  Condensation is getting past unit and into lines going to instruments.  Unit needs to be removed and sent for repairs.  At 1520 GMT, shut down compressor and removed unit.  Whatman model FR-IR 75-45-12VDC, SN 7545-0236.  RTS at 1600 GMT.

3/24/11

CAM

Removed hatch assembly due to it being in the field of view of SAS instruments.   Stored hatch assembly under table in Optical trailer.

Instrument is no longer in service

5/19/09

PAD

Remotely accessed computer at 1205 GMT and Scheduler Vi program is not running.  Started Scheduler Vi.  After start up, status shows:  Waiting till start time.  At 1428 GMT, status shows:  Environment check complete.  Bad Weather.  Filled LN2 dewers and selected "Force Good Weather" box.  Instrument started taking measurements.  RTS at 1443 GMT.

5/18/09

PAD

Message shown:  Quitting will abort all running Vis.  Cleared message.  Status shows:  Environment check complete.  Bad Weather.  Weather is good outside.  Checked "Force Good Weather" box and system started taking measurements at 1404 GMT.  At 1516 GMT, status shows:  Closing OPUS.  Exited and restarted Scheduler Vi program.  Status shows:  Bad Weather.  Selected "Force Good Weather" again and system started taking measurements.  RTS at 1531 GMT.

5/15/09

PAD

At 1305 GMT, Status shows:  Environment check complete-Bad Weather.  Error message is shown:  Quitting will abort all running Vis.  At 1305 GM, cleared error message.  Selected "Force Good Weather" box and instrument started taking measurements.  At 1355 GMT, status shows:  Closing OPUS.  Rebooted computer and selected "Force Good Weather" box again.  At 1815 GMT, unchecked "Force Good Weather" box due to rain possible later today.  RTS at 1830 GMT.

5/13/09

PAD

At 1503 GMT, Status shows:  Environment check complete-Bad Weather.  Selected "Force Good Weather" box and instrument started taking measurements.  At 1516 GMT, status shows:  Closing OPUS.  Exited/ restarted program and selected "Force Good Weather" box again.  At 2058 GMT, unchecked "Force Good Weather" box due to rain possible later today.  RTS at 2113 GMT.

4/8/09

PAD

Entered Optical building and heard air leak from dry air system for SWS and Sorti.  Found leak was o-ring on drain valve of inline filter.  Replaced o-ring and leak was stopped.  Found another leak in tubing between regulator and filter.  R/R with copper tubing.  RTS at 1545 GMT.

3/19/09

PAD

Status shows:  Closing OPUS.  Tracker is rotating.  Overcast sky.  Closed Scheduler Vi program and disconnected power to tracker.  At 1451 GMT, skies began to clear.  Connected power to tracker and started Scheduler Vi program.  At 1607 GMT, remotely accessed computer and message was shown.  Quitting will abort all running Vi's.  Canceled error message. 

3/18/09

PAD

Scheduler Vi program is not running.  Hatch is open and tracker is not rotating.  Started Scheduler Vi program.  Status shows:  Waiting till start time.  At 1520 message is shown:  Quitting will abort all running Vi's.  Hatch was still open and not rotating.  Rebooted computer and power cycled system.  Tracker started rotating while computer was rebooting.  Status shows:  Waiting till start time.  At 1728 GMT, remotely access computer and error was shown:  AQP Read Time Out.  Selected Retry for error.  No data files are shown.  Status shows:  Taking measurement 1 of 2.  Checked on instrument remotely later and program is running with data files being produced. 

3/17/09

PAD

Status shows:  Closing OPUS.  Tracker is rotating.  Started scheduler Vi program.  Status shows:  Waiting till start time.  At 1613 GMT, message shows:  Quitting will abort all running Vi's.  No data files are shown.  Canceled message.  At 1852 GMT, program was running and had data files from 1104-1237 MST. 

3/16/09

PAD

At 1852 GMT, observed message on screen:  Quitting will abort all running Vi's.  Canceled message.  Status shows:  Waiting specified time till next measurements.  There were data files from 1104 to 1237 MST.

3/16/09

PAD

Mentor has made changes to operating software and requests instrument started and ran for a period of time to analyze collected data.  Connected power to tracker and rebooted computer at 1357 GMT.  Labview program started and status shows:  Waiting till start time.  Tracker started rotating after power was restored and will occasionally lock onto sun.  Will continue to monitor as time allows today. 

2/12/09

PAD

Status shows:  Starting OPUS, but nothing is updating.  Restarted program and error was displayed:  Error 7 occurred at Open File+.vi.  Cleared error and computer rebooted on its own.  Program started and status shows:  Waiting on scheduled time.

3/12/08

PAD

No air pressure at inline dryer supplying SORTI and SWS instruments.  Reset breaker for air compressor in optical trailer.  Air compressor started and pressure stabilized.  RTS at 1543 GMT.

10/29/07

PAD

SORTI tracker has not been operating since 10/25/07.  At 1910 GMT, Mark Klassen and I T/S tracker problem.  T/S found 28V DC at connecter located in tracker hatch. Tracker turns to limit counterclockwise then stops and does not rotate clockwise.  Notified instrument mentor.  Waiting on instructions.  RTS at 2030 GMT.

10/3/07

CAM

The site lost power for an extended time beginning around 1940 GMT on 10/2 during a severe thunderstorm. Instrument ran for a short time on UPS power and then shut down.  Power was restored after normal business hours yesterday.  Computer was operational this morning but indicated a Data Acq Error.  Selected retry on the error message window and system RTS status O.K. at 1406 GMT.

7/31/07

MK

The program was not running. Restarted software, system is operational.

7/17/07

PAD

Two error messages are displayed:  Quitting will abort all running Vi's and  Bruker FTIR.  AQP Read Timeout.   Canceled error messages but display keeps showing Bruker FTIR error message.  At 1459 GMT stopped program and shut down computer and instrument.  Powered up instrument and then computer.  Error message came up but did not return after canceling.  RTS at 1514 GMT.

7/16/07

PAD

Two error messages are displayed:  Quitting will abort all running Vi's and  Bruker FTIR.  AQP Read Timeout.   Canceled error messages but display keeps showing Bruker FTIR error message.  After consulting with mentor, stopped program and shut down computer and instrument.  Powered up instrument and then computer.  Error message came up but did not return after canceling.  RTS at 1445 GMT.

6/14/07

PAD

Inlet pressure on dryer controller for SWS and SORTI shows 0 psi.  Reset breaker for air compressor.  RTS at 1445 GMT.

6/13/07

PAD

Inlet pressure on dryer controller for SWS and SORTI shows 60 psi.  Reset breaker for air compressor.  RTS at 1508 GMT.

5/29/07

MK

The power at the CF had failed some time over the weekend causing the UPS to shutdown due to low voltage on batteries. Powered on UPS at 1434 GMT, restarted SORTI computer and software at 1440 GMT.

4/17/07

PAD

No programs are running on computer.  Restarted Scheduler Vi program.  RTS at 1425 GMT.

4/9/07

PAD

Computer is active but no program is running.  Restarted Scheduler Vi. program at 1404 GMT.  RTS at 1419 GMT.

4/3/07

MK

The SORTI software is not running at this time. Restarted software at 1526 GMT. RTS at 1541 GMT.

4/2/07

PAD

Log file for 3/31/07 at 2019 shows:  Measurement successful.  Closing OPUS.  Restarted Scheduler Vi. program at 1612 GMT.  RTS at 1627 GMT.

3/20/07

PAD

Storms at CF tripped UPS located in Optical trailer.  Last successful collection was at 0300 GMT.  Following instruments were affected:  AERI01, MWR, RSS, SWS, TSI-880, SORTI01, and VCEIL25K.  Facilities reset UPS at 1600 GMT.  Restarted computers that did not restart after power was restored.  RTS at 1635 GMT.

3/14/07

PAD

Log entry for 3/13/07 at 2159 GMT shows:  Measurement successful.  Closing OPUS.  Restarted Scheduler Vi program.  Program restarted correctly.  RTS at 1436 GMT.

3/13/07

PAD

Log entry for 3/12/07 at 2107 GMT shows:  Measurement successful.  Closing OPUS.  Restarted Scheduler Vi program.  Program restarted correctly.  RTS at 1502 GMT.

3/12/07

PAD

Last log entry for 3/10/07 at 1646 GMT shows:  Measurement successful.  Closing OPUS.  Restarted Scheduler Vi program.  Program restarted correctly.  RTS at 1441 GMT.

3/7/07

MK

The SORTI software was observed locked up, restarted program at 1648 GMT.

3/5/07

MK

The SORTI software was observed locked up, restarted program at 1640 GMT.

3/1/07

CAM

Remote status check at 1800 GMT found the SORTI program had shut down again.  Scheduler program status showed Closing Opus and Opus program was in fact closed.  Log file showed last entry of Closing Opus at 1553 GMT.  Placing the instrument on UPS power seems to have not solved the problem.  Shut down and rebooted computer at 1803 GMT.  Program restarted and RTS at 1807 GMT.    Also noticed some possible problems with the time zone settings and how the clock is set.  The computer time is accurate, but does not agree with the modified dates of the log files.  Will discuss this with mentor.

2/28/07

CAM

The SORTI instrument is not connected to UPS power.  Suspect this may be the cause of frequent program hang-ups due to the instrument resetting during power bumps.  Discussed this with instrument mentor and he agreed the instrument should be connected to UPS power.    Went to shut the instrument down and found the program was hung up again.  Scheduler program showed Closing Opus.  The last 2 entries in the log file indicated successful measurement 1 of 1 at 1738 GMT and then closing Opus at 1739 GMT.  Removed the instrument from raw power and plugged into UPS power at 2005 GMT.  Restarted scheduler program and RTS at 2008 GMT.

2/27/07

CAM

While adjusting the SORTI alignment and using the measurement program in Opus, the site experienced a hard power bump.   Observed that the program froze up and the instrument appeared to do some sort of reset.  A mirror on a belt system slid over into the light path between the input mirror and first internal mirror, thus blocking the light from entering the main optical body.  Had to shut down and restart the program to get the SORTI to restart and the measurement program running again.  Observed that the power supply for the instrument is not on UPS power.  Suspect power bumps may be the cause of frequent program hang ups that have been occurring for the past several months.  Notified mentor of the observations. 

2/27/07

CAM

Instrument mentor suggested Site Ops try moving the SORTI instrument slightly on the table to improve alignment and signal counts.  Stopped the program at 1817 GMT.  Recorded max count with 3.XPM filter as 3380 and 7.XPM filter as 4430.  Marked the current instrument location on the table and then tried rotating it slightly to improve the light angle on the input mirror and increase the count.  No matter where the instrument was located, the maximum count obtained always remained in the range of 3200-3400.  Ended up with the SORTI located on the original marks made by the mentor the last time the instrument had been removed for maintenance.  Maximum count with 3.XPM filter at that point was 3260.  Noticed that the instrument did not appear to be level after having moved it.  The NE leg under the main instrument body was not solid on the table.  Screwed the adjustable leg out to level it and noticed the count would increase as that corner of the instrument was raised.  Adjusted the leg so that it was solid on the table and the count ended up at 3690.  Notified mentor of the results and requested instructions for leveling the instrument.  Restarted program and RTS at 2058 GMT.

2/26/07

CAM

Instrument mentor requested site ops check the SORTI alignment due to lower than expected signals.  Stopped the program at 1825 GMT.  Loaded the appropriate file/filter for alignment.  The amplitude was approximately 3400.  Adjusting the input mirror only increased the amplitude slightly to around 3600.  It appears as though the instrument itself is rotated slightly from where it should be and thus causing the light to enter at an angle.  The light coming from the input mirror is not passing cleanly through the instrument cover, but hits the cover on the right hand side.  However, on the first internal mirror, the light is hitting too far to the left.  Took pictures and sent an email to mentor describing what was observed.  T/S will continue.  Restarted the program and RTS at 1945 GMT.

2/26/07

PAD

Log file for 2/23/07 at 2004 GMT shows:  Measurement successful.  Closing OPUS.  Rebooted computer and instrument restarted correctly.  RTS at 1525 GMT.

2/23/07

PAD

Log file shows:  Successful measurement.  Closing OPUS.  Restarted Scheduler. Vi program.  RTS at 1601 GMT.

2/22/07

PAD

Log file for 2/21/07 at 2047 shows:  Successful measurement.  Closing OPUS.  Restarted Scheduler. Vi program.  Program restarted correctly.  RTS at 1549 GMT.

2/20/07

PAD

Log file for 2/17/07 at 1555 shows:  Successful measurement.  Closing OPUS.  Restarted Scheduler. Vi program.  Program restarted correctly.  RTS at 1605 GMT.

2/15/07

PAD

Log file shows:  Closing OPUS at 2104 GMT on 2/14/07.  Restarted program.  Instrument status shows:  Environment check complete.  Bad weather.  RTS at 1540 GMT.

2/5/07

PAD

Log file for 2/2/27 at 1707 GMT shows:  Measurement failed.  Closing OPUS.  Rebooted computer at 1534 GMT.  Program appears to have restarted correctly.  RTS at 1549 GMT.

1/30/07

PAD

Log file for 1/29/07 at 1858 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1518 GMT.  RTS at 1533 GMT.

1/29/07

PAD

Log file for 1/26/07 at 2107 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1536 GMT.  RTS at 1551 GMT.

1/26/07

PAD

Log file for 1/25/07 at 2055 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1548 GMT.  RTS at 1603 GMT.

1/25/07

PAD

Log file on 1/24/07 at 1654 shows:  initializing.  Labview is up, but nothing is running.  Closed out program and restarted.  Error message shown:  Error- 200324 occurred at DAXmx write (Digital 1D Bool 1Chan 1Samp). vi.  Instrument status shows:  Starting OPUS.  Nothing appears to be running at this time.  Rebooted computer and restarted program.  Program restarted correctly.  RTS at 1738 GMT.

1/24/07

PAD

Log file on 1/23/07 at 1628 shows:  initializing.  Labview is up, but nothing is running.  Closed out program and restarted.  Error message shown:  Error- 200324 occurred at DAXmx write (Digital 1D Bool 1Chan 1Samp). vi.  Instrument status shows:  Starting OPUS.  Nothing appears to be running at this time.  Will check tomorrow.  RTS at 1728 GMT.

1/22/07

PAD

Log file for 1/19/07 at 1922 GMT shows:  Measurement failed.  Closing OPUS.  Restarted Labview program.  RTS at 1611 GMT.

1/19/07

PAD

Log file for 1/18/07 at 1959 GMT shows:  Measurement failed.  Closing OPUS.  Restarted LabView program at 1533 GMT.  RTS at 1548 GMT.

1/18/07

PAD

Log file for 1/17/07 at 2102 GMT shows:  Measurement failed.  Closing OPUS.  Restarted Lab view program at 1538 GMT.  RTS at 1553 GMT.

1/17/07

PAD

Log file for 1/16/07 at 1756 shows:  Measurement failed.  Closing OPUS.  Restarted Lab view program at 1530 GMT.  RTS at 1545 GMT.

1/16/07

PAD

Instrument was not restarted on 1/12/07 due to approaching ice storm.  Instrument program needs to be restarted.  Restarted Lab view program at 1553 GMT.  Hatch opened correctly and instrument started taking measurements.  RTS at 1608 GMT.

1/11/07

PAD

Tower hatch is closed.  Log file for 011007 at 1830 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1521 GMT.  RTS at 1536 GMT. 

1/10/07

PAD

Tower hatch is closed.  Log file for 010907 at 1733 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1628 GMT.  RTS at 1643 GMT. 

1/9/07

PAD

Tower hatch is closed.  Log file for 010807 at 1744 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1522 GMT.    RTS at 1537 GMT.

1/8/07

PAD

Tower hatch is closed.  Log file for 010607 at 1840 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program at 1545 GMT.  Instrument started taking measurements.    RTS at 1600 GMT. 

1/4/07

PAD

Last entry in log file for 1/3/07 at 1739 GMT shows:  Measurement failed.  Closing OPUS.  Program is not running and hatch is not open.  Restarted program at 1506 GMT.  RTS at 1521 GMT.

1/3/07

PAD

Last entry in log file for 1/2/07 at 2017 GMT shows:  Measurement failed.  Closing OPUS.  Program is not running and hatch is not open.  Restarted program at 1531 GMT.  RTS at 1546 GMT.

1/2/07

PAD

Instrument is not active.  Last log file entry for 1/1/07 at 1646 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program.  RTS at 1524 GMT.

12/27/06

PAD

Log file for 12/26/06 shows:  Closing OPUS at 1745 GMT.  Program is not running.  Restarted Sorti program at 2055 GMT.  RTS at 2110 GMT.

12/26/06

PAD

Log file for 12/22/06 shows:  Closing OPUS at 1957 GMT.  Program is not running.  Restarted Sorti program at 1537 GMT.  RTS at 1552 GMT.

12/18/06

PAD

Last log entry on 12/15/06 at 1737 GMT:  Measurement failed.  Closing OPUS.  Restarted program.  Instrument status shows:  Environment check complete.  Bad weather.  RTS at 1600 GMT.

12/15/06

PAD

Last log entry on 12/14/06 at 1803 GMT:  Measurement failed.  Closing OPUS.  Restarted program.  Instrument status shows:  Environment check complete.  Bad weather.

12/13/06

PAD

Instrument status shows:  Closing OPUS.  Nothing is updating on weather page.  Last log file entry for 12/12/06 shows:  Measurement failed.  Closing OPUS.  Restarted program.  Instrument status shows:  Environment check complete.  Bad weather.

12/12/06

PAD

Instrument status shows:  Closing OPUS.  Nothing is updating on weather page.  Last log file entry for 12/11/06 at 1844 GMT shows:  Measurement failed.  Closing OPUS.  Restarted program.  Labview. exe error message was displayed followed by Error-50150 occurred at DAQmxCreate Channel (DO-Digital Output) vi.  error message.  Printed screen of error messages and saved on computer desktop.  Stopped and restarted program again.  Instrument status shows:  Environment check complete.  Bad weather. 

12/11/06

PAD

Instrument status shows:  Closing OPUS.  Nothing is updating on weather page.  Last log file entry for 12/8/06 shows:  Measurement failed.  Closing OPUS.  Stopped and restarted program.  Instrument status shows:  Environment check complete.  Bad weather. 

12/7/06

PAD

Instrument status shows:  Closing OPUS.  Most recent log file was on 12/7/06 at 1810 GMT:  Measurement failed.  Closing OPUS.  Restarted program at 1510 GMT.  Instrument status shows:  Environment check complete.  Bad weather.  RTS at 1525 GMT.

12/5/06

PAD

Instrument status shows:  Closing OPUS.  Most recent log file was on 12/4/06 at 1932 GMT:  Measurement failed.  Closing OPUS.   Restarted program at 1609 GMT.  Instrument status show:  Environment check complete.  Bad weather.  RTS at 1615 GMT.

12/4/06

PAD

Instrument status shows: Closing OPUS.  Most recent log file was on 12/1/06.  Checked sunseeker tower hatch to make sure no ice had built up on it from recent freezing rain and snow before restarting program.  Restarted program at 1552 GMT.  Instrument status show:  Environment check complete.  Bad weather.  RTS at 1607 GMT.

11/28/06

PAD

At 1525 GMT, observed sunseeker tower turning back and forth but hatch is closed.  Instrument status shows:  Environment check complete... Bad weather.  The sun is shining at this time at the CF.  I checked the force good weather box on the scheduler. vi page.  Hatch opened and instrument started taking measurements.  Unchecked force good weather box at 1619 GMT.  Did not want to leave instrument running in force good weather mode as there is a winter storm predicted at the CF tomorrow.  RTS at 1634 GMT.                       

11/27/06

PAD

Instrument status shows:  Closing OPUS.  Weather graphs are not updating.  Looked in log files.  Last entry was on 11/24/06 at 1646:39- Measurement failed... Closing OPUS.  Restarted Labview program.  Notified mentor of ongoing problem with instrument stopping measurements.  Troubleshooting at this time.  Overcast at CF today.  RTS at 1532 GMT.

11/24/06

PAD

Labview program shows instrument status:  Closing OPUS.  Graphs on weather page are not updating.  Looked at log files on C drive and last entry for 11/22/06 at 2145 GMT reads:  Measurement failed/ Closing OPUS.   Restarted program and Instrument status shows:  Environment check complete...Bad Weather. Scattered clouds at this time at CF. I later observed sunseeker tower was open and instrument appeared to be locked onto sun.  RTS at 1506 GMT.

11/22/06

PAD

At 2013 GMT, Labview window is minimized and will not open.  Window open labeled:  Please vind the VI named "Scheduler.vi"  Log file shows measurement failed/ closing OPUS at 18:05:36 GMT.  Restarted Labview program.  Program appears to have restarted correctly.  RTS at 2028 GMT.

11/22/06

PAD

At 1545 GMT, instrument status says:  Closing OPUS.  No graphs are updating on weather screen.  Restarted Labview program but instrument status says:  Environment check complete- Bad weather.  It is clear at the CF.  Called instrument mentor.  Mentor request file name changed in C: data/ vi.  Changed file name Bruker.lib to Bruker-11-20-2006.lib.  Changed Bruker-11-09-2006.lib to Brukerlib.  Remote Admin. shows someone is logged on with remote screen 130.253.240.131 Mentor is not logged on at this time.  Stopped and restarted remote server terminating connection.  RTS at 1650 GMT.

11/15/06

PAD

Instrument mentor request Sorti computer clock changed from CST to GMT.  Reset computer clock to GMT time at 1636 GMT.  RTS at 1651 GMT.      

11/2/06

Mentor

New instrument mentor (Tom Hawat) on site and reinstalled the SORTI. 

7/12/04

MR

The instrument mentor wants the instrument shut down and stored in place.  Shut the instrument down at 2015 GMT.  Checked the hatch and the hatch was closed.  Turned off all power supplies to the instrument.  Left the dry air supply running so the optics will stay dry.  RTS at 2030 GMT.

7/12/04

MR

The instrument mentor has returned the SORTI hard drive and wants it installed and tested.  Installed the hard drive and powered the system up.  Will allow the system to run for approximately 1 hour for testing and then will shut the instrument down.  RTS at 1630 GMT.

6/16/04

PAD

Instrument mentor requested hard drive removed from SORTI computer and sent to him for data retrieval.   Shut down SORTI program, computer and instrument at 1433 GMT.  Tim Grove removed Western Digital hard drive model # AC28400-00RT, SN WM628  029 8990.  Hard drive has been sent to mentor.  RTS at 1448 GMT.

5/20/04

PAD

Instrument mentor requested files be moved from E drive to D drive to free up memory on E.  Tim Grove stopped program on 5/19/04 at 2030 and started transferring data.  Restarted program on 5/19/04 at 1407 GMT.  Both D and F drives are full.  Instrument mentor has been notified.  RTS at 1422 GMT.

4/16/04

TMG

The SORTI computer E: drive has only 32KB of available free space.  Mentor requested some data be moved from E: to F: to free more space. Shutdown the SORTI program @ 1633 GMT and moved Data from E:\OBS\ to F:\ drive. There is 2,080,704 KB Free Space on E:\ Drive now.  Placed the following data folders on F: drive. These folders are from 20030802, 20030806 thru 20030825, and 20031217.  Rebooted the computer at 1727 GMT.

2/18/04

PAD

At 1540 during PM rounds, hatch was observed closed and instrument was rotating.  At 1624 GMT, stopped program and rebooted computer.  Notified instrument mentor.  RTS at 1639 GMT.

2/12/04

PAD

During PM rounds, SORTI hatch was observed cycling between open and closed.  Instrument would lock on sun and hatch would close, then hatch would open and instrument would seek and lock on sun.  Stopped program at 1643 GMT.  Rebooted computer and power cycled silver rack below instrument.  Displayed showed a Master. CMD message:  Waiting to run task d:\ roams\ dowtask2\ asam.tsk .35 .32 90 at 12:29:59.  Checked instrument later.  Instrument seemed to operate normal.  RTS at 1658 GMT.

2/3/04

PAD

On 2/2/04, hatch was observed frozen shut due to rain and cold temperatures.  Stopped program at 1523 GMT.    On 2/3/04 at 1906 GMT, checked hatch to make sure ice had melted.  Powered computer and made sure hatch operated properly.  RTS at 1921 GMT.

2/2/04

PAD

Recent rain followed by snow and lower temperatures has frozen hatch closed.  Stopped program and shut down computer at 1523 GMT.

1/20/04

PAD

During PM rounds, inlet pressure on regulator for purge generator was 22 psi.   Outlet flow pressure was 0.  Checked breakers for compressor.  Reset breaker #6 in breaker box at 1608 GMT.  Gauge on compressor showed 150 psi.  At 1625 GMT, regulator inlet pressure slowly built up to 60 psi.  Purge generator cycled and inlet pressure increased to 100 psi.  Outlet flow pressure rose to 10 psi.  Notified facilities department and Dan Nelson said he would have someone check compressor for ice in line and drain moisture from compressor if necessary.  RTS at 1625 GMT.

1/15/04

PAD

Instrument hatch was open but instrument was not tracking sun.  Computer was up and display screen looked normal.  Rebooted computer at 1541 GMT.  Made sure program started correctly.  Shut hatch to make sure computer was communicating with instrument.  Computer displayed Master CMD message:  Master control waiting to run task d:\ roams dowtask 2 asamrun. tsk .35 .05  90 at 11:30:20.  Checked instrument at 11:30 CST.  Master CMD showed same message with time changed to 12:29:59.  Contacted instrument mentor.  Waiting on reply.

8/7-8/8/03

Mentor

Instrument mentor on site to repair system.  SORTI brought back on line 8/8/03.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

12/11/02

JEM

The mentor has requested that the SORTI computer be shipped to him so he can make the necessary repairs.  Disconnected all cables from the computer @1520 GMT and took the computer to George so it could be shipped out for repair.  RTS @1535 GMT.

12/4/02

JEM

Due to the problem with the SORTI computer software, the SORTI needs to be shutdown until the mentor can visit the site to correct the problem.  Powered off the computer @1905 GMT.  Powered the instrument off @1907 GMT.  Instrument will remain offline until necessary repairs are made.  RTS @1915 GMT.

12/3/02

JEM

The SORTI is still down awaiting repairs to the OS/2 Operating System.

12/02/02

JEM

The SORTI computer has locked up and is not responding to any commands.  Power cycled the SORTI computer @1535 GMT.  When the computer came back online @1540 GMT, the following error message was observed.  "Your desktop could not be found in the OS2.ini file.  A temporary desktop will be created."  I notified the computer department of this problem.  RTS @1550 GMT.

10/4/02

JEM

The following error was observed.  OPUS:Error in opening file.  E:\OBS\20021004\A9CA0407.0.  Clicked cancel on the error message.  Hard drive E: remains full.  Will contact the mentor again about this to inform him the hard drive still needs to have space freed up on it.    RTS @1436 GMT.

9/25/02

JEM

The following error message was observed.  OPUS:Error in opening file.  E:\OBS\20020925\A9C92507.0.  Clicked cancel on the error message.  Status Normal.  Notified the mentor about these frequent errors.  RTS @1524 GMT.

9/23/02

JEM

The following error message was observed on the monitor:  OPUS:Error in opening file.  E:\OBS\20020923\A9C92307.0  This error has occurred frequently.  The only exception is the date is always the current date.  Clicked cancel.  Status normal.  Will notify mentor directly of the errors.  RTS @1448 GMT.

9/16/02

JEM

The following error message was observed.  OPUS: Error in opening file

E:\OBS\20020916\A9C91607.0.  There is an option to retry or cancel.  Retry brings the same error back up. Clicked cancel.  Status returned to normal.  RTS @1515 GMT.

9/9/02

JEM

The SORTI computer was found locked up.  Power cycled the SORTI computer @1500 GMT.  After the computer boot status returned to normal.  RTS @1514 GMT.

8/30/02

MR

The computer was locked up.  Rebooted the computer.  RTS at 1625 GMT.

8/28/02

MR

The computer was locked up.  Rebooted the computer at 1533 GMT.  RTS at 1548 GMT.

8/22/02

KT

There was no response from the computer.  The computer appeared to be locked up.  At 1623 GMT rebooted the computer.  At 1628 GMT the computer came on line and the program appeared to be operating.  RTS at 1635 GMT. 

8/20/02

MR

The computer was locked up.  Rebooted the computer.  System came back up at 1430 GMT.  RTS at 1443 GMT.

8/12/02

KT

The computer program appeared to be locked up.  At 1629 GMT rebooted the computer.  At 1632 GMT the computer came on line, but the mouse was not operating.  At 1637 GMT rebooted the computer once again.  After second reboot the computer came on line and the mouse control was back.  RTS at 1640 GMT. 

8/12/02

KT

The hatch was found open and the instrument was continuously rotating.  At 1626 GMT power cycled the instrument.  After power cycling the instrument, the rotating stopped but the hatch was still open.  At 1630 GMT selected the hatch close icon on the computer.  At 1635 GMT hatch finally closed.  RTS at 1640 GMT. 

7/29/02

MR

The computer is locked up.  Rebooted the computer at 1509 GMT.  RTS at 1524 GMT.

7/19/02

CT

A Acquisition Processor error had occurred. The message stated: Acquisition Processor!! Error 2: Laser A Missing!!  Should this message be disabled in future attempts?  YES   NO   At 1418 the NO option was selected. The error message was cleared. The mentor was notified by email of the problem. Status OK. RTS at 1433 GMT. 

7/17/02

CT

The error message that was displayed was " Acquisition processor!!  Error 3 : laser B missing!!" Should this message be disabled in future attempts? YES or NO The mentor was contacted by Mike Rainwater. At the mentor's request the NO option was selected at 1727 GMT. The instrument continued with the sampling. Status OK. The mentor was contacted concerning the outcome of the action taken. RTS at 1742 GMT.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

2/5/02

DN

Due to an extended power outage, the IM has requested that we provide this instrument with a source of dry nitrogen.  Installed nitrogen cylinder with regulator and flow meter.  Adjusted flow rate at 1 SCFH.

11/19/01

Mentor

Mentor powered down instrument on 11/18/01.  Mentor performing troubleshooting.  Instrument is down until further notice.

11/02/01

CJ

Found program locked up at 1705 GMT. Rebooted computer at 1705 GMT. RTS at 1720 GMT

11/01/2001

CJ

Program found locked up at 1454 GMT. Rebooted computer at 1454 GMT. RTS at 1509 GMT.

10/31/01

CJ

Found the program locked up at 1523 GMT.  Rebooted computer at 1523 GMT. RTS at 1538 GMT.

10/30/01

CJ

Program found locked up at 1520 GMT.  Rebooted computer at 1520 GMT. RTS at 1535 GMT.

10/29/01

CJ

Found program terminated at 1657 GMT. Restarted program at 1657 GMT. RTS at 1712 GMT

10/18/01

CJ

Found program locked up at 1446 GMT. Rebooted computer at 1449 GMT. RTS at 1501 GMT.

10/15/2001

CJ

Computer display was locked up. Rebooted at 1630. RTS at 1645.

10/12/01

MR

The computer was locked up. Rebooted the computer at 1804 GMT.  RTS at 1819.

10/2/01

SS

Computer needed to be rebooted.  Computer was rebooted at 1420 GMT.  When the reboot was complete, the data returned to normal.  RTS 1435 GMT.  Status was OK at this time.

9/27/01

KT

The computer was found to be locked up.  At 1833 GMT rebooted the computer.  RTS at 1848 GMT. 

9/13/01

KT

The instrument was just rotating back and forth.  At 1355 GMT rebooted the computer.  At1359 GMT the computer came on line and the instrument stopped rotating and began normal operation.  RTS at 1408 GMT. 

8/20/01

MR

The power was lost to the site on 8/15/01 from 0030 GMT to 0400 GMT.   This instrument was found off. Powered on the computer.  System came up and started working at 1416 GMT.  RTS at 1428 GMT.

8/14/01

JEM

At 1905 GMT noticed an error message on SORTI computer that states; No spectrum selected; could not start file:  Rebooted computer at 1905 GMT.  Normal data and operations restored at 1920 GMT.

8/9/01

JEM

At 1415 GMT SORTI computer display was showing an error 'disk is full.’ At this time a 'soft' reboot was performed.  Computer program was operational at 1430 GMT.

8/2/01

MR

The DIO control board for the SORTI came in and needs to be installed. Installed the DIO board int he SORTI computer.  Filled the instrument with LN2 and notified the mentor.  He wants the instrument to run and will notify us when to restart the daily LN2 fills.

7/18/01

MR

The instrument mentor called and wants to try and power up the SORTI and run some tests on it. Powered up the SORTI and started the tests.  During one of the tests the SORTI hatch was supposed to open but would not.  After several attempts the mentor wanted me to check to see if the system was getting power and that the power supply was working.  Checked for input power and it checked good.  The internal power supply on the SORTI hatch controller was not working.  The mentor wants to check some drawings and said he would call me back on 7/19/01 to continue with the testing and T/S.  Powered the system back down and will await further instructions from the mentor.

7/11/01

MR

The instrument mentor called today and wants to power up and test the system. Powered up the system.  The system came up and the AQP board started but then the system indicated a lost link.  T/S over the phone with Tom Stephens found a bad power supply module.  Removed the module and returned it to Bruker industries for repair.  Conner Flynn will send us a replacement unit.  RTS at 1900 GMT.

7/06/01

MR

The AQP board has come in and needs to be reinstalled. Reinstalled the AQP board.  Will wait to talk to the mentor to power up the system and run the tests on it.  Notified the mentor.  RTS at 1700.

5/9/01

MR

The instrument mentor wants to run the instrument at 1500 for the IOP. Called the instrument mentor and notified him that the instrument was ready to run. He had me turn the instrument on but the AQP board would not boot up. Switched the board around and put it in another bus slot but this did not cure the problem. The mentor wants to send another board out by FedEx so it could be installed tomorrow.

4/17/01

MK

The SORTI hatch was sent to mentor for repairs.

4/16/01

MK

The SORTI hatch was found blown off and laying on the ground. Shut the system down and notified mentor.

4/13/01

MR

The computer was locked up. Rebooted the computer at 1445 GMT. RTS at 1500 GMT.

3/22/01

MR

The computer is locked up. Rebooted the system at 1534 GMT. RTS @1549.

3/14/01

MR

The computer was locked up. Rebooted the computer at 1546 GMT. RTS at 1601 GMT.

3/12/01

MR

The computer is locked up. Rebooted the computer at 1528 GMT. RTS at 1543 GMT.

3/8/01

MR

The computer was locked up. Rebooted the computer at 1700 GMT.

3/6/01

MR

The computer was locked up. Rebooted the computer at 1612 GMT. RTS at 1628 GMT.

3/5/01

MR

The computer has locked up. Rebooted the computer at 1537 GMT. RTS at 1552.

2/19/01

MR

The mentor is onsite performing the following maintenance. 1: Checked status of LN2 Detector hold time - > 24 hours OK. 2: Cleaned electrical contacts in CPU and instrument - Corrosion problems were evident in the operations of the instrument. The controller board AQP was not booting properly. Checks out OK. 3: Checked signal levels - within specs. 4: Measured interior mirror for replacement. 5: Adjusted scanner cable in the FTS. 6: Configured Power to FTS so that it is always on. 7: Replace reed relay for FTS scanner control. 8: Cleaned up data directories. 9: Performed PM on sunseeker assembly. This included: cleaning contacts on PC boards and interconnects, cleaning cable contacts, cleaned tachometer windings and brushes, adjusted elevation feedback potentiometer, fixed shutter closure so that the shutter is captured by the nylon guard and removed 3 mud dobber nests in tracker. 10: Backed up software. 11: Ran tests to check performance and stability.

2/9/01

MR

The computer was locked up. Rebooted the computer at 1533 GMT. RTS at 1548 GMT.

1/31/01

MR

The computer was locked up. Rebooted the computer at 1550 GMT. RTS at 1605 GMT.

1/30/01

MR

The computer was locked up. Rebooted the computer at 1506 GMT. RTS at 1521 GMT.

1/22/01

MR

The computer was locked up. Rebooted the computer at 1550 GMT. RTS at 1605 GMT.

1/16/01

MR

The computer was locked up. Rebooted the computer at 1642. RTS at 1657 GMT.

1/12/01

MR

The computer was locked up. Rebooted the computer at 1458 GMT. RTS at 1515 GMT.

1/8/01

MR

The computer display was locked up. Rebooted the computer at 1540 GMT. RTS at 1555 GMT.

1/5/01

MR

The computer was locked up. Rebooted the computer at 1600 GMT. RTS at 1615 GMT.

1/2/01

MR

The computer was locked up. Rebooted the computer at 1730 GMT. RTS at 1745 GMT.

12/29/00

KT

The computer was locked up and needed to be rebooted. At 1623 GMT rebooted the computer. At 1625 GMT the computer came on line. RTS at 1635 GMT.

12/12/00

MR

The air compressor was noticed to be running without shutting off. Observed that the safety valve on top of the compressor was jarred slightly open and not allowing the compressor to build up enough pressure to shut down. Reseated the valve and the leak stopped. The compressor stopped shortly after this.

12/8/00

MR

The computer is locked up. Rebooted the computer at 1532. RTS at 1546.

12/07/00

KT

The computer had locked up and the instrument was rotating from side to side. At 1529 GMT performed reboot of the computer. At 1533 GMT start up was complete and the master control waiting to run task window was up. The instrument had ceased its rotation. RTS at 1541 GMT.

12/5/00

MR

The computer has locked up. Rebooted the computer at 1600 GMT. RTS at 1615 GMT.

12/4/00

MR

The computer was locked up. Rebooted the computer at 1530. RTS at 1545.

11/28/00

MR

The program was locked up. Rebooted the computer. The program came back up and started running at 1523 GMT. RTS at 1535 GMT.

11/21/00

MR

The computer was locked up. Rebooted the computer at 1516 GMT. RTS at 1531 GMT.

11/20/00

MR

The computer was locked up. Rebooted the computer at 1542. System started to run at 1545. RTS at 1557.

11/16/00

MR

The computer was locked up and would not respond to keyboard commands. Rebooted the computer at 1545 GMT. System came up and restarted at 1547 GMT. Status O.K.

11/14/00

MR

The computer was locked up. Rebooted the computer at 1638 GMT. System came back up at 1640 GMT. RTS at 1653.

11/13/00

KT

The computer was found locked up during morning rounds. At 1620 GMT rebooted the computer. At 1622 GMT the computer came on line and Master control waiting to run task appeared. RTS at 1635 GMT.

11/09/00

MR

The computer would not respond to the keyboard or the mouse. Rebooted the computer at 1541 GMT. RTS at 1555 GMT.

11/3/00

MR

The computer was found locked up. Rebooted the computer at 1512 GMT. RTS at 1527 GMT.

10/30/00

MR

The computer was locked up. Rebooted the computer at 1534 GMT. System was RTS at 1549 GMT.

10/24/00

JEM

The SORTI computer was found to be locked up. Tried rebooting the computer using CTRL-ALT-DEL, but the computer was unresponsive. Had to turn the computer power off and back on again with the power button. The SORTI appeared to be operational again @1457 GMT. RTS @1457 GMT.

10/20/00

MR

The computer is locked up and will not respond. Rebooted the computer at 1424 GMT. System came back up and started running at 149 GMT. RTS at 1441 GMT.

10/18/00

JEM

The SORTI computer was found to be locked up @1358 GMT. Rebooted the computer @1358 GMT. The computer came back online and was operating properly @1410 GMT. RTS @1413 GMT.

10/17/00

JEM

The SORTI computer was found to be locked up @1325 GMT. Rebooted the computer @1326 GMT. The computer came back online and was operating properly @1340 GMT. RTS @1345 GMT.

10/11/00

MR

The computer was locked up. Rebooted at 1430 GMT.

10/5/00

MR

The computer was locked up. Rebooted the computer at 1345 GMT. System started to run at 1400 GMT.

9/25/00

MR

The SORTI computer was locked up with no response from the keyboard. Rebooted the computer at 1430 GMT. RTS at 1445.

9/21/00

MR

Rebooted the computer at 1453 GMT. RTS at 1503 GMT. Computer was locked up.

9/19/2000

MR

There was no response from the keyboard or the mouse. Power cycled the computer at 1745. System came back up at 1747. Status O.K.

9/18/2000

MR

Computer was locked up and would not respond to keyboard. Rebooted the computer at 1735 GMT. System restarted at 1740. RTS at 1750.

8/4/2000

MR

Only the desktop was observed. The program looked as if it was not running. Rebooted the computer. The computer came back up and looked the same way. It was discovered that this was normal and the status program was run and it showed that the instrument was working. RTS at 1448.

5/30/00

ZW

The SORTI computer was locked up, with no keyboard, or mouse control. At 1410 GMT using the (Ctrl-Alt-Del) rebooted the SORTI computer. Reboot was successful. Status OK. RTS 1430 GMT.

5/24/00

MR

Power to the optical trailer is to be shut down for ½ the day. Gracefully shut down the computer at 1456. Powered back up at 2001.

5/22/00

KT

At 1410 GMT observed that the computer was locked up. At 1412 GMT rebooted computer. RTS at 1425 GMT.

5/15/00

KT

At 1730 GMT mentor request reboot of instrument computer. At 1733 GMT computer was rebooted. RTS at 1745 GMT.

5/10/00

KT

At 1854 GMT the computer needed to be rebooted. At 1855 GMT computer came on line. RTS at 1919 GMT.

5/3/00

KT

Mentor request reboot of instrument computer. At 1841 GMT rebooted computer. RTS at 1856 GMT.

5/2/00

KT

At 1805 GMT mentor request reboot of the instrument computer. Rebooted computer. RTS at 1820 GMT.

5/1/00

KT

Mentor emailed request to reboot instrument computer. At 1725 GMT computer was rebooted. RTS at 1740 GMT.

4/25/00

KT

At 1846 GMT the instrument computer needed to be rebooted. At 1847 GMT rebooted computer. RTS at 1902 GMT.

4/24/00

KT

Mentor requested reboot of instrument computer. At 1749 GMT rebooted computer. RTS at 1806 GMT.

4/21/00

KT

Observed that the computer was locked up. At 1743 GMT rebooted the computer. RTS at 1758 GMT.

4/20/00

KT

At 1753 GMT the computer needed to be rebooted. At 1755 GMT rebooted computer. RTS at 1810 GMT.

4/17/00

KT

At 1756 GMT mentor emailed request to reboot the sorti computer. At 1758 GMT rebooted computer. RTS at 1811 GMT.

4/10/00

KT

At 1800 GMT mentor emailed request to reboot instrument computer. At 1803 rebooted computer. RTS at 1815 GMT.

4/4/00

KT

Mentor emailed request to reboot SORTI computer. At 1734 GMT rebooted computer. RTS at 1736 GMT.

4/3/00

KT

The computer was found locked up. At 1800 GMT rebooted the computer. At 1805 GMT computer came on line with Master control waiting to run task. RTS at 1806 GMT.

3/29/00

KT

Mentor request that the instrument computer be rebooted. At 1923 GMT rebooted the computer. At 1925 GMT computer came on line with Master control waiting to run task. RTS at 1928 GMT.

3/21/00

KT

At 1432 GMT observed the same window DOVTASK.CMD appeared. At 1433 GMT rebooted the SORTI computer. At 1434 GMT the computer came on line with Master Control waiting to run task. RTS at 1437 GMT.

3/20/00

KT

At 1536 GMT there was a window labeled DOVTASK.CMD that appeared on the screen and also appeared that the computer had locked up. At 1541 GMT rebooted SORTI computer. At 1543 GMT the computer came on line with Master control waiting to run task. RTS at 1546 GMT.

3/9/00

KT

At 1950 GMT the SORTI computer appeared to be locked up. At 1951 GMT rebooted the computer. RTS at 1954 GMT.

3/1/00

KT

At 2008 GMT mentor requested reboot of the computer. RTS at 2012 GMT.

2/29/00

KT

At 1929 GMT mentor request reboot of computer. At 1930 GMT rebooted computer. At 1934 GMT computer came on line Master control waiting to run task. RTS at 1935 GMT.

2/23/00

KT

At 2040 GMT mentor request reboot of the SORTI computer. At 2043 GMT rebooted computer. RTS at 2047 GMT.

2/16/00

KT

At 1920 GMT there was an error message stating AQP (port:336) not found. At 1921 GMT rebooted computer. At 1924 GMT system came on line with Master control waiting to run task. RTS at 1925 GMT.

2/15/00

KT

At 1845 GMT there was an error message stating AQP (port:336) not found. At 1847 GMT rebooted computer. At 1848 GMT system came on line with Master control waiting to run task. RTS at 1850 GMT.

2/11/00

KT

There was an error message at 1445 GMT stating AQP (port:336) not found. At 1447 GMT rebooted computer. At 1448 GMT computer came online with Master control waiting to run task. RTS at 1450 GMT.

2/1/00

KT

Mentor needed computer rebooted. At 1836 GMT rebooted computer. At 1838 GMT system came on line. RTS at 1839 GMT.

1/25/00

KT

At 1840 GMT the computer needed to be rebooted per email request. RTS at 1845 GMT.

1/21/00

KT

The SORTI computer had error AQP (port:336) not found. At 1834 GMT rebooted computer. At 1836 GMT computer came on line. RTS at 1837 GMT.

1/14/00

KT

Mentor needed the system rebooted. At 1930 GMT rebooted computer. At 1936 GMT system came on line. RTS at 1940 GMT.

1/13/00

KT

At 1926 GMT the instrument was found rotating with the hatch closed. There was also an error message IFS120M-AQP(port:336) not found. At 1928 GMT rebooted system. RTS at 1932 GMT.

1/13/2000

CAM

The SORTI computer appeared to be locked up. A box in the lower right hand corner of the screen read Booting AQP. A window was open in the center of the screen with an error message which read AQP Not Found. Rebooted the computer at 1545 GMT. RTS at 1555 GMT.

1/12/2000

CAM

Instrument mentor notified Site Ops that he was unable to ping or telnet into the SORTI computer remotely. Mentor requested that the computer be rebooted. Site Ops could successfully ping the SORTI computer across the network, and could also ping other computers outside the local network. Rebooted the computer as mentor requested at 1715 GMT. RTS at 1720 GMT.

1/11/2000

KT

Mentor needed instrument rebooted. At 1927 GMT computer was rebooted. RTS at 1940 GMT.

1/10/2000

KT

Mentor was unable to communicate with instrument. At 1854 GMT rebooted computer. RTS at 1904 GMT.

1/6/2000

CM

While retiling the Optical trailer floor, contractors plugged a compressor into a UPS outlet which caused the breaker to trip. The UPS and all Optical trailer instruments lost power and had to be restarted. Removed the compressor from the UPS outlet and reset the breaker. Restarted all Optical trailer instruments. RTS status O.K.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will need to be shut down. Shutdown Sorti at 1435 GMT, restarted Sorti at 1615 GMT. RTS at 1615 GMT.

11/8/99

KT

At 1920 GMT the Purge Flow rate value was reading 12. Value needed to be adjusted back to 10. At 1923 GMT adjusted value to 10. RTS at 1925 GMT.

10/15/99

KT

At 1858 GMT the outside compressor was powered off due to leaking connecting putty from an elbow connection on the side of the compressor.

8/9/99

MR

Rebooted the computer due to site power construction. RTS at 1755.

8/3/99

MR

Blew the dust off the mirror using compressed air. RTS at 1445.

6/15/99

MR

The purge gas generator air moisture indicator is yellow. Talked to Dan Nelson and he has some on order. Dan suggested putting a small amount of desiccant in the bottom of the filter housing below the filter to be able to monitor if the air is dry. Placed small amount of desiccant in the bottom of the housing. RTS at 1645.

6/14/99

CAM

The SORTI was taken off line at 1850 GMT on 6/2/99 pending installation of the new air compressor. The compressor installation is now complete and SORTI needs to be restarted. Placed the beam splitter back in the SORTI and started the program at 1550 GMT. RTS status O.K.

6/2/99

CAM

The air compressor used to purge the SORTI has become too unreliable for continuous operation. The compressor frequently trips the circuit breaker and shuts down. This could result in costly damage to the SORTI beam splitter. Selected killops on the SORTI PC to shut the system down at 1850 GMT. Removed the beam splitter from the SORTI and placed in the desiccant tank. Shut the compressor breaker off and unplugged compressor at 1855 GMT. Notified mentor of the actions taken. SORTI will remain off line until the new compressor is installed, likely within a week.

6/1/99

CAM

The beam splitter was removed for the weekend on Friday at mentors request. Reinstalled the beam splitter at 1335 GMT. Selected 'Reset Ops' on the computer to restart the program. RTS status O.K. at 1340 GMT.

5/28/99

MR

Activated the killops program. Removed the mirror at 2115.

5/24/99

CAM

The beam splitter was removed for the weekend on Friday at mentors request. Reinstalled the beam splitter at 1345 GMT. Selected 'Reset Ops' on the computer to restart the program. RTS status O.K. at 1350 GMT.

5/21/99

CAM

Selected 'Kill OPS' command to shut the system down at 2100 GMT. Removed the beam splitter and stored in the desiccant tank. This will be done each Friday at mentors request and then reinstalled on Monday morning to prevent high humidity damage to mirror. Once the new compressor is operational this process will no longer be required.

5/19/99

MR

Reset the breaker to the air compresser. RTS at 1350.

5/18/99

Mentor

Tom Stephen on site to reinstall the SORTI computer. System back on line.

4/7/9

CAM

Instrument is off line per mentor. The computer has been removed for repair.

1/5/99

MW

Had to shut the system down to remove the failed sun tracker assembly. Shut down the computer at 1745 GMT. Powered all components off and placed the beam-splitter in the desiccant tank. Removed the tracker and motor from the tower. Covered the opening at the top of the tower with a tarp and secured the tarp to the tower with some rope. Will ship the tracker back to the mentor for repair.

12/28/98

MW

The SORTI sun seeker hatch was open this morning and the seeker was rotating back and forth. The internal mirror was rotating up and down, making a loud thumping noise as it hit the bottom end of its rotation. The SORTI computer was locked up and would not respond to operator commands. The time indicated on the display was 02:02:12 PM. Rebooted the computer at 1556 GMT. Attempted to synchronize the sun seeker by using the 'sun_on' and 'sun_off' commands but the hatch would not respond. The hatch motor sounded like it was rotating, but the hatch itself was not moving. Called the instrument mentor and left a message. The instrument mentor called back and said he would disable the sun seeker and hatch assembly remotely. He also indicated that he will generate a work request at a later time stating what action needs to be taken next.

12/22/98

MW

The normal SORTI ‘Master.cmd’ display window was not present and it appeared that the SORTI was not functioning. Rebooted the computer but the 'Master.cmd' window did not appear. Called the instrument mentor and he was able to correct the problem remotely. RTS at 2040 GMT.

12/15/98

MW

The SORTI01 computer was locked up and would not accept operator commands. The time the system crashed according to the display was 01:08:41 PM local time. The sunseeker hatch was in the open position. Rebooted the computer at 1832 GMT. Noticed a message on the pop-up window after the reboot which indicated that it could not write a file to "E:\OBS" because the disk was full. Verified that there was no available space on "E:\" drive. Synchronized the sunseeker hatch. Rebooted again at 1839 GMT. The system locked up again at 1851 GMT after the next scan because the OPUS program could not open a necessary file in the "E:\OBS" directory. Rebooted again at 1856 GMT. Closed the SORTI program. Closed the sunseeker hatch. Called the instrument mentor and left a message.

12/9/98

MW

The SORTI computer was locked up and the sunseeker hatch was in the open position. The time indicated on the display was 0816 local time. The computer would not respond to operator commands. Rebooted the computer at 1628 GMT. Synchronized the sunseeker hatch and rebooted again at 1635 GMT. RTS at 1637 GMT.

12/7/98

KT

At the time of inspection the SORTI instrument was trying to seek the sun and the computer had locked up. At 1530 GMT rebooted the computer which reset the SORTI back to the home position but did not close the hatch. At 1540 GMT after talking with Chris Martin we were instructed to engage the limit switch on the instrument to the closed position and type in sun_on on the computer. Then, after disengaging the limit switch we typed in sun_off which in turn placed the SORTI instrument back in the home position and also closed the hatch. RTS at 1655 GMT.

11/24/98

MW

The SORTI computer was locked up this morning. The time the system locked up was 08:32:27 AM according to the display. Called the instrument mentor and informed him of this ongoing problem. He was aware of the problem and asked that the system be rebooted. Rebooted the computer at 1610 GMT. Synchronized the sun seeker hatch position and rebooted again at 1618 GMT. Returned to service at 1820 GMT.

11/23/98

SS

Daily checks revealed that the computer is locked up. The computer is locked up with 08:30:07 AM displayed. Rebooted the computer at 1630 with Control+Alt.+Del. command. The sun tracker is now out of synchronization. Synchronized the sun tracker. Left a voice mail for Tom Stephens detailing actions. RTS at 1645 GMT.

11/20/98

SS

Daily checks revealed that the computer locked up at 08:27. Rebooted the computer using Control+Alt.+Del. at 1631 GMT. The sun tracker is open and tracking not synchronized with the instrument. Synchronized the sun tracker with the instrument. Left a voice mail for Tom Stephens.

11/19/98

MW

The SORTI sunseeker hatch was open in the 'sun off' position. Rain drops were observed on the seeker mirror. 1659 GMT - Opened an OS/2 window and changed directory to 'D:\OK_OPS' and typed 'SUN_ON'. The seeker began rotating and the hatch closed. Typed 'SUN_ON' once again and the hatch opened. Typed 'SUN_OFF' and the hatch closed and the seeker stopped rotating. Shut down the operating system and rebooted at 1703 GMT. The system came back online at 1705 GMT.

11/16/98

SS

The computer is locked up and displays a time of 07:52:59 AM. Rebooted with Control+Alt.+Del at 1606 GMT. the computer came up and looks normal. The sun seeker tower is out of synchronization with the instrument. Synchronized the sun seeker tower with the instrument at 1619 GMT. RTS at 1620 GMT.

11/13/98

SS

Daily check revealed that the computer was locked up and the sun seeker was searching. Rebooted the computer at 1518 GMT. The sun seeker is out of synchronization with the instrument. Synchronized the sun seeker with the instrument. RTS at 1535 GMT.

11/12/98

SS

Daily checks revealed that the Sorti computer is locked up and the sun seeker tower is open. Rebooted the computer at 1635. The sun seeker was open and out of synchronization with the instrument. Synchronized the sun seeker by commanding the sun seeker on and manually tripping the micro switches to allow the sun seeker to remain open. Commanded the sun seeker off to get it into synchronization with the instrument. RTS at 1640 GMT.

10/30/98

MK

The SORTI computer was locked up and the sunseeker was searching for the sun. Rebooted computer at 1606 GMT. Synchronized the sunseeker tower with the computer program. RTS at 1620 GMT.

10/9/98

ZW

SORTI is down at this time the Mentor knows this.

9/24/98

KT

System is still powered on. Still awaiting reply from mentor.

9/21/98

KT

System is powered on but there is no data scrolling that is visible. Called and left message with mentor and awaiting reply.

9/14/98

SS

The computer is locked up and will not respond to either mouse or key board controls. The monitor shows 06:31 PM and Master Control waiting to run Task - d:\roams\htch_cls at 23:49. The sun seeker hatch is open but not tracking the sun. Rebooted the computer at 1400 GMT with the Control+Alt+Delete command. The computer came up but the dun seeker tower was not synchronized with the instrument. Synchronized the sun seeker tower with the instrument. RTS at 1415 GMT.

9/10/98

SS

Daily checks revealed that the computer locked up at 06:37 PM. Computer would not respond to keyboard or mouse commands. Rebooted the computer using Control+Alt+delete command at 1420 GMT. Computer came up at 1422 GMT with Master Control waiting to run Task - d:\roams\dowtask2 asamrun.tsk at 08:43. Synchronized the sun seeker tower with the instrument. RTS at 1430 GMT.

9/8/98

SS

The computer was locked up time reads 06:43 PM and the screen reads Master Control waiting to run Task - d:\roams\htch_cls at 23:49. The sun seeker is closed. Rebooted the computer at 1518 GMT using the Control+alt+delete command. And the computer came up with the clock reading 09:22 AM and running and the screen reading Master Control waiting to run Task - d:\roams\dowtask2 asamrun.tsk at 12:29:59. RTS at 1530 GMT.

9/4/98

SS

Daily checks revealed that the computer locked up at 06:45 AM and will not recognize mouse of key board commands. Rebooted the computer at 1330 GMT with the Control+Alt+Delete command. Computer then came up with: Master Control waiting to run Task -d:\roams\dowtask2 asamrun.tsk at 08:38:19. The sun seeker tower is open and out of synchronization with the instrument. Commanded sun on and tripped the micro switch so that the hatch would stay open and be synchronized with the sun seeker then commanded sun off to synchronize the sun seeker with the instrument. RTS at 1345 GMT.

9/3/98

SS

Daily check revealed the computer to be locked up at 06:47 PM. Display indicates Master control waiting to run Task - d:\roams\htch_cls at 23:49:59. Rebooted the computer using the control+alt+delete command. During reboot the computer comes up with the error device on line 118 not found and is ignored and line 104 is ignored press enter to continue. The computer does come up after pressing enter and appears to be normal. The sun seeker tower appears to be in synchronization with the instrument. Called Tom Stephens and left message describing problem.

9/2/98

SS

Daily checks revealed that the computer stopped at 1:40 PM MST and will not respond to mouse or keyboard controls. Rebooted the computer using Control+Alt+Delete command. Computer came up and appears to be normal but the sun seeker tower is now out of synchronization with the instrument. Commanded the sun seeker off with no change. Commanded sun seeker on and tripped the micro switch for the cover to stop the cover from closing. Commanded the sun seeker off and the sun seeker returned to the proper condition for the next sampling period. RTS at 1500 GMT

9/1/98

SS

The computer is locked up with "OPUS:Error in opening file E:\OBS\19980901A9890108.0". The sun seeker tower is open. Clicked on Cancel and the Program closed the sun seeker tower then went into a waiting to run task mode. Computer appears to be normal now the sun seeker tower and instrument program are synchronized. RTS at 1500 GMT.

8/31/98

SS

Daily check revealed that the computer was locked up and would not respond to mouse commands clock read 1:40 PM. Rebooted the computer with Control+Alt+Delete command. Computer came up properly but now the sun seeker is open and out of synchronization with the instrument. Commanded the sun seeker on and tripped the micro switch to allow the cover to remain open. Then commanded the sun seeker off and the sun seeker returned to the proper position to be synchronized with the instrument.

8/28/98

SS

The computer is locked up and will not recognize mouse commands. Rebooted the computer using the control+alt+delete command at 1500 GMT. Computer locked up with line 104 of the config.sys not recognized presses enter and computer came up. RTS at 1515 GMT.

8/27/98

SS

The computer clock is locked up and it will not respond to mouse commands. Used control+alt+delete to reboot computer then performed a computer shut down and power off before rebooting computer again. Found the sun seeker cover to be out of synchronization with the sun seeker program. Secured the micro switch for the open position and performed sun_on. This synchronized the cover and program then performed sun_off to synchronize the sun seeker with the instrument. System now appears to be ready and waiting for the next collection period. RTS at 1500 GMT.

8/24/98

SS

The SORTI computer clock reads 7:09PM and is not responding to keyboard or mouse commands. Rebooted computer at 1526 GMT.

8/14/98

ZW

The SORTI is down at this time. The mentor is on site, and working on this instrument.

8/10/98

MK

During morning PM checks, noticed that the SORTI computer was locked up and the tracker was in seek mode with the hatch open. The mirror had dried water spots on it due to rain overnight. Rebooted the computer at 1455 GMT. Closed the hatch with the "sun_off" command. Rebooted again at 1515 GMT. RTS at 1530 GMT.

7/29/98

MW

The SORTI computer was locked up and the tracker was in seek mode with the hatch open this morning. The mirror has rain drops on it due to rain overnight. Rebooted the computer at 1435 GMT. Closed the hatch with the "sun_off" command. Rebooted again at 1446 GMT. RTS at 1450 GMT.

7/23/98

MW

Site ops has procured a new generator to supply air to the SORTI. The air compressor was installed and the SORTI was restarted at 2015 GMT. Filled with LN2 at 2035 GMT. RTS at 2045 GMT.

7/1/98

MW

The SORTI was shut down today due to failure of the air compressor which feeds air to the instrument. Shut the system down at 1320 GMT and removed the KBr beam splitter and placed it in the desiccator. The system will remain down until the air compressor is back online.

7/1/98

MW

The air compressor which feeds air to the SORTI has malfunctioned and is not running. The compressor motor tripped its motor circuit breaker. When the motor circuit breaker was reset, the motor tripped the AC power circuit breaker. After the AC power breaker was reset, the motor would not activate. T/S in progress. Informed the SORTI mentor of the problem.

6/30/98

MK

The SORTI computer was locked up and the sunseeker is open and searching for the sun while it was raining. Rebooted the SORTI computer at 1437 GMT. Accessed the D: drive and changed to the ok_ops directory and ran the hatch close program at 1445 GMT. The hatch in now closed and the computer status is normal.

6/15/98

SS

Daily check revealed SORTI computer locked up at 01:04:08 PM and the sun seeker continuously searching. Computer does not respond to keyboard or mouse. Power cycled computer at 1603 GMT. After the computer came up the sun seeker was not searching but the cover was open. Opened an OS2 window on the computer and changed to D:\ok_ops\ and entered sun_on. The sun seeker began searching but the hatch closed. Entered sun_off and the sun seeker quit searching but the hatch opened. Entered sun_on, taped the open micro switch closed, and then entered sun_off. Now the sun seeker hatch is synced with the seeker. RTS at 1630 GMT.

6/4/98

CAM

The SORTI sun seeker hatch was open and continuously rotating with overcast conditions. The computer appeared to be locked up. The cursor would move but the mouse buttons had no effect. Notified Tom Stephen who advised to reboot computer. Rebooted computer at 1450 GMT. Program came back on line at 1455 GMT. The sun seeker hatch stopped rotating but remained open. Mentor advised he would close the hatch remotely. Checked hatch status at 1555 and found it was still open. Mentor indicated he was unable to access the system remotely. Closed hatch locally at 1600 GMT as instructed by mentor. After the hatch closed the system hung up again with the mouse having no effect. Rebooted computer at 1601 GMT. The mouse driver failed to load on boot up. Hard booted computer at 1604 GMT. RTS status O.K. at 1607 GMT.

5/26/98

MK

Due to a power failure on 5/25/98 at 0425 GMT, the SORTI computer has locked up. Shutdown and restarted SORTI computer at 1445 GMT. status OK. RTS at 1515 GMT.

5/15/98

SS

Due to power failure restarted instrument at 1525 GMT.

4/20/98

MW

The SORTI sunseeker hatch was found open while it was raining. The mirror was not moving and the sunseeker itself was not rotating. It appeared to be in the sun_off position with the hatch open. Gave a 'sun_off' command at 1905 GMT. Nothing happened. Issued a 'sun_on' command and the mirror began moving, the sunseeker began rotating, and the hatch closed. Issued a 'sun_off' command and depressed the hatch limit switch to stop the hatch from opening fully. Held the limit switch and issued a sun_on command, the hatch opened, the mirror started moving and the sunseeker began rotating. Somehow (possibly by a power outage) the hatch was not synchronized with the sunseeker. When the sunseeker wished to go to sun_on mode, the hatch would close. When the instrument wished to go to the sun_off mode, the hatch opened. Had to 'trick' the instrument by depressing the hatch limit switches and typing the right commands. Rebooted the system at 1920 GMT. RTS.

4/2/908

CAM

Tom Stephen phoned and notified Site Ops that the SORTI computer had locked up with the hatch in an undetermined state. Due to possible thunderstorms overnight the computer needed to be rebooted and the hatch verified as closed. Rebooted SORTI computer at 2220 GMT. Tom remotely closed the hatch and system RTS at 2230 GMT.

4/1/98

MW

The SORTI computer was locked up this morning and the hatch assembly was rotating back and forth from one limit switch to another. Could not get any response from the system. Rebooted the computer at 1520 GMT. The system came back online at 1525 GMT. The instrument mentor called after the reboot and requested that the SORTI be completely shut down and cold booted since he could not access the system remotely. Backed out of all running applications and cold booted the computer at 1735 GMT. The system came back online at 1740 GMT.

3/31/98

ZW

The SORTI display was locked up. Rebooted computer at 1600 GMT. Display returned to normal. RTS 1615 GMT.

3/9/98

MW

The SORTI hatch was opening and closing continuously, preventing the sun tracker from staying locked onto the sun. The computer locked up at 1645 GMT and would not respond to the keyboard or mouse. The mentor asked for the hatch to be closed using the "HTCH_CLS" command until he could troubleshoot the problem further. Power cycled the computer at 2155 GMT. Closed the hatch at 2212 GMT. The next day (3/10/98), the instrument mentor restarted the system remotely at 1840 GMT. RTS at 1845 GMT.

3/6/98

TG

Mentor requested that remote communications be enabled on the SORTI computer. Technician found computer locked up at 1545 GMT. Rebooted computer at 1557 GMT. Technician configured TCP/IP software to Auto-start Inetd, telnetd, and ftpd applications. Reboot computer at 1605 GMT and Tom Stephens logged into computer. System RTS at 1610 GMT.

3/2/98

Mentor

Mentor arrived on site 3/2 for maintenance to get the system operational. SORTI operational on 3/3/98.

12/18/97

MW

The SORTI01 sun tracker is not operational. Instrument mentor requested that the hatch be closed and power disconnected. Also, the mentor requested that the SORTI computer be shut down. The hatch closed on its own at 1620 GMT after being open for at least 45 minutes. Shut down the SORTI computer and disconnected the three conductor power connector which feeds the tracker assembly at 1622 GMT.

12/17/97

MW

The SORTI sun tracker is not tracking the sun. The hatch was open, the mirror was tilting up and down, but the tracker was facing to the West and was not rotating East-West. Disconnected the tracker from the azimuth motor assembly and verified both cable connections to the azimuth motor. Both cables were connected. Replaced the tracker on the azimuth motor and reconnected power. The tracker instantly moved all the way to the West, past the limit switch, up against the rubber stop bumper. Manually rotated the tracker all the way to the East to the Eastern limit switch. When the tracker was released, it immediately rotated all the way past its Western limit switch at a high rate of speed and hit the rubber stop bumper. Removed the tracker from the azimuth motor and lowered the screw on the motor assembly which depresses the Western limit switch to ensure that the switch was being depressed all the way. Replaced the tracker on the azimuth motor and reapplied power. Once again, the tracker instantly rotated past its Western limit switch up against the rubber stop bumper. Suspect the motor control circuitry is faulty. Informed the SORTI sun tracker mentor of the problem. T/S still in progress.

12/10/97

MW

The SORTI has been down since 10/6/97 while the sun tracker was being repaired. Reinstalled the SORTI tracker on 12/10/97 at 2000 GMT. Rebooted the SORTI computer at 2010 GMT. RTS at 2015 GMT.

12/8/97

CAM

The computer was locked up and indicated that the device driver on line 88 of config.sys file was not installed. Pressed enter to continue boot up process. The system however remains off line awaiting shipment of sun seeker mounting hardware from mentor.

11/12/97

CAM

Pressed the space bar and 'enter' at 1535 GMT to unlock screen saver before noticing that the monitor was not powered on. Powered the monitor on and the system was rebooting. Apparently the computer had been locked up and pressing 'enter' allowed the system to reboot. SORTI remains off line however due to sun seeker cover being removed for repairs.

10/6/97

MW

The sun seeker hatch is in the open position but the seeker is not pointing towards the sun. Normally, the hatch is open only at the top of the hour for about 15 minutes (if conditions permit). The hatch should be closed at this time. Rebooted the SORTI computer at 1457 GMT. No effect. Checked all AC power connections. Everything appears to be plugged in. Contacted instrument mentor at 1505 GMT and left message.

10/6/97

MW

The SORTI hatch is not closing and the tracker is not moving as of 10/6/97. On 10/6/97, the voltage to the sun seeker hatch was measured at 28 V from the output jack of the power supply at 2100 GMT. There is a fault somewhere inside the sun seeker assembly. The hatch assembly was covered with plastic and taped up at 2120 GMT (10/6/97) to keep rain out of the unit until further troubleshooting can be performed.

10/6/97

MW

The SORTI computer was locked up this morning and emitting a continuous rapid beeping noise (approximately 5 beeps per second.) It appeared to be in DOS mode as if it had rebooted. There was an error message indicating a problem with the mouse.sys driver in the config.sys file and another message at the bottom prompting the user to press 'enter' to continue. Pressed 'enter' at 1404 GMT. Shortly afterwards the beeping noise stopped but the screen went black and the system did not continue to boot up as it should. Pressed 'ctrl-alt-del' at 1406 GMT to see if the computer would soft boot. The system stopped booting after the BIOS read and the beeping noise resumed. Power cycled the computer at 1408 GMT. This time the computer began booting normally but the beeping noise returned. Hit 'enter' during the boot up process and the beeping noise halted. The system came all the way up into OS/2 and the SORTI program started. RTS at 1409 GMT.

10/1/97

MW

The SORTI was not running this morning. The display indicated that the system was in DOS mode and there was a message at the bottom of the screen prompting the user to press enter to continue. Pressed enter on the keyboard at 1557 GMT. The system appears to have come back on-line.

7/28/97

CM/CW

AC power to the optical trailer had to be shut off at 1845 GMT for construction. Performed graceful shut down at 1840 GMT. AC power restored at 1905 GMT and system RTS status O.K. at 1910 GMT.

3/26/97

MLS

The display screen on the SORTI has been password-locked. I will try to get in contact with Tom Stephens sometime today to find out what the password is so the techs can do a meaningful observation of the systems performance.

3/20/97

ZW/CM

Daily PM the screen saver on the computer was locked up called Tom Stephen. Tom asked me to power cycle the computer. Power cycled computer at 1620 GMT. Tom is still troubleshooting this problem.

9/11/96

Mentor

On site to * R/R system. Testing incomplete and left off line until replacement parts can be installed. * old - 00 * new - 01

7/3/96

CAM

Daily check found display with "Acquisition processor" error, Laser A "off", scanner LED flashing, ready LED not lit at 1505 GMT. Notified mentor who instructed to leave off line until he can visit to make repairs. System will be off line.

6/13/96

CAM

Daily check RH = 44%. R/R desiccant at 1400 GMT.

6/12/96

CAM

Daily check found display " insufficient space on target area" error. Notified mentor at 1330 GMT

6/12/96

CAM

Cleared error and RTS at 2045 GMT. RH = 44%. R/R desiccant at 2115 GMT but RH is still high. Checked all seals and found one loose. Repaired and RTS at 2130 GMT.

5/29/96

TMG

Daily check found display "OPUS" error opening file. Disk full C:\OPUS\Measa 9652912.0 selected. Performed system shut down and notified mentor at 2145 GMT.

5/28/96

TMG

Daily check RH = 38.4%. R/R desiccant at 1349 GMT.

5/17/96

CAM

Daily check RH = 34.4%. R/R desiccant at 1408 GMT.

5/13/96

CAM

Daily check RH = 32%. R/R desiccant with EBBR type at 2000 GMT.

5/8/96

CAM

Daily check RH = 46%. R/R desiccant at 1345 GMT.

5/6/96

CAM

Daily check RH = 48%. R/R desiccant at 1400 GMT.

2/13/96

TMG/GRS

AC power secured for construction at 1520 GMT. RTS at 1946 GMT.

1/29/96

TMG/CAM

AC power secured for construction at 1424 GMT. RTS at 2044 GMT.

 12/27/95

TMG/CAM

AC power secured for construction at 1500 GMT.